A closed system initially containing 1.000 × 10-3 M H2 and 2.000 × 10-3 M I2 at 448℃ is allowed to reach equilibrium. At equilibrium, the HI concentration is 1.87 × 10-3 M. Given that the reaction equation is H2(g) + I2(g) ↔ 2 HI(g), find:
a. Equilibrium concentrations of the two reactants
b. Kc of the reaction taking place.
1
Expert's answer
2020-10-29T05:24:06-0400

Solution:

Chemical reaction:


H2(g) + I2(g) ↔ 2HI(g)


Given:

[H]in = 10-3M

[I]in = 2*10-3M

[HI]eq = 1.87*10-3M


this problem is solved by using ICE table method:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 H_2 & I_2 & 2HI \\ \hline
 1*10^{-3}M & 2*10^{-3}M & 0 \\
 \hdashline
 x & x & 2x\\ \hdashline [H_{2}]_{eq} & [I_{2}]_{eq} & 1.87*10^{-3}M
\end{array}"


Firstly we should find x(change) in the ICE table:


2x=1.87*10-3

x=9.35*10-4


a) Equilibrium concentrations of the two reactants:


[H]eq = 10-3 - 9.35*10-4 = 6.5 * 10-5 M

[I]eq = 2*10-3 - 9.35*10-4 = 1.065 * 10-3 M


b) Kc of the reaction taking place:


"K_c= \\dfrac{[HI]^2}{[H][I]}\t= \\dfrac{(1.87*10^{-3})^2}{(1.065*10^{-3})(6.5*10^{-5})}\t=50.5"


"K_c=50.5"